# From a Tiny Slip to a Giant Leap: An LLM-Based Simulation for Fake News Evolution

### v1

@@ 3.3. Propagation Role-Aware Agent
(2) $f_{role}=f_{r_{i}}(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}).$
@@
For the role in our model, spreaders may combine and amplify sensational aspects of the news, commentators may add personal interpretations or opinions, verifiers may cross-reference information before accepting or sharing it, and bystanders typically retain their previous news content $S_{i}^{t}=S_{i}^{t-1}$ unless significantly influenced (Sun et al., 2023). The prompt for the role-specific reintroduction function $f_{r_{i}}$ is in Appendix A.
3.3.3. Memory and Reflection.
@@
In our simulation, agents engage with their neighbors each day, leading to an updated version of the news. However, owing to the potentially vast volume of interactions, storing all of them in detail is impractical. Therefore, agents are equipped with a hierarchical memory system comprising short-term memory (STM) $M_{i}^{S}$ and long-term memory (LTM) $M_{i}^{L}.$ The STM stores recent interactions and immediate information received, while the LTM accumulates knowledge over time, influencing future information processing and news updates. After interactions, agents reflect and reintroduce the news through a memory consolidation function:
(3) $M_{i}^{L,t}=g(f_{L}(M_{i}^{L,t-1}),f_{S}(M_{i}^{S,t})),$
@@
where $g(\cdot)$ integrates new information into LTM, allowing agents to exhibit dynamic behaviors such as gradually changing their opinion on a topic or reinforcing existing opinions. The function prompt for short-term memory $f_{S}$ and long-term memory $f_{L}$ is in Appendix A.
3.3.4. Decision-Making Process.
@@
In our FUSE framework, each agent’s opinion evolves through a reasoning process influenced by their role, persona, and interactions with others. After daily interactions and memory updates, agents reflect on the news content they hold, leading to gradual changes in their opinions. This process integrates the agent’s role in propagation—whether they are a spreader, commentator, verifier, or bystander—with their cognitive functions. The decision-making process for agent $a_{i}$ at time $t$ is modeled as:
(4) $S_{i}^{t}=f_{dm}(S_{i}^{t-1},m_{i}^{L,t-1},r_{i},\mathcal{P}_{i}).$
@@
The function prompt $f_{dm}$ is in Appendix A, which captures how agents integrate new information with their existing opinions, taking into account their role in the decision-making process. For example, the reasoning of spreaders may lead to greater changes in $S_{i}^{t}$. Commentators often add subjective nuances to the news content through their reasoning. Verifiers are aiming to correct inaccuracies and potentially reduce misinformation. Bystanders have minimal engagement; their reasoning leads to minor or no changes in their held content.
@@
3.4. News Evolution Simulator
The News Evolution Simulator (NES) serves as the core environment where news content propagates and evolves over time through interactions among agents. This module simulates the flow of information within a social network structure $\mathcal{G}=(\mathcal{A},\mathcal{E})$, enabling the study of how true news gradually transforms into fake news due to the influence of agent behaviors, memory systems, and social interactions.
NES models a dynamic, evolving social network where agents interact based on predefined topologies. Specifically, the framework explores network structures such as random networks, where the edges $\mathcal{E}$ between agents $a_{i}\in\mathcal{A}$ are formed randomly. This simulates environments where individuals are not tightly clustered. In scale-free networks, certain agents, or hubs, have significantly more edges, simulating influential actors or ”super-spreaders” in social media. High-clustering networks are modeled to form communities where agents are highly connected within their local group, reflecting real-world interaction patterns seen in tightly-knit social circles. The network structure $\mathcal{G}$ dictates which agents interact daily, influencing the rate and manner in which news spreads and evolves in the system. The overall algorithm is shown in Algorithm 1.
3.4.1. Intervention Mechanisms
A critical feature of the NES is its ability to simulate interventions aimed at countering the evolution of fake news, thereby contributing to the development of a more responsible web environment. When the degree of deviation between the current news content $S_{i}^{t}$ and the original news $S_{0}$ surpasses a predefined threshold, an official agent is introduced into the system to correct misinformation. This official agent represents authoritative sources that provide verified information, to mitigate the spread and impact of distorted or false narratives.
The intervention process starts with continuous monitoring of the deviation between the news content held by each agent and the original news. Once the deviation exceeds a critical threshold, the official agent is triggered to take action. This agent issues official announcements based on reliable sources, targeting agents most likely to propagate or exacerbate misinformation. The prompt for the official agent is in Appendix A.
4. FUSE-EVAL: News Evolution Analysis
To systematically measure how true news evolves into fake news within our simulation, we propose a comprehensive evaluation framework named FUSE-EVAL. This framework consists of two sets of metrics: Content Deviation Metrics and Statistical Deviation Metrics, which together provide a detailed understanding of how misinformation evolves within the simulated environment.
4.1. Content Deviation Metrics
The Content Deviation Metrics assess the deviation of the news content across multiple dimensions by quantifying changes in specific aspects of the news. FUSE-EVAL evaluates the news content based on six key dimensions. (1) Sentiment Shift (SS) measures the change in emotional tone between the original news content and its evolved version (Lu et al., 2022; Ma et al., 2021). Sentiment plays a crucial role in how information is perceived and shared; shifts in sentiment can indicate the introduction of bias or emotional manipulation. (2) New Information Introduced (NII) assesses the extent to which additional information, not present in the original news, has been incorporated during the evolution process (Arif et al., 2016; Wang et al., 2017). The introduction of new facts, details, or claims can significantly alter the original message, potentially leading to misinformation if the new information is inaccurate or misleading. (3) Certainty Shift (CS) evaluates changes in the level of confidence or assertiveness expressed in the news content (Krafft and Spiro, 2019; Kim and Yoon, 2022). Shifts from definitive statements to speculative language (or vice versa) can influence the perceived credibility and urgency of the information. (4) Stylistic Shift (STS) examines changes in writing style, tone, and linguistic features (Wu et al., 2024; Alsaif and Aldossari, 2023). Alterations in style—such as moving from formal to informal language, changing sentence complexity, or using different rhetorical devices—can affect readability and audience engagement. (5) Temporal Shift (TS) measures changes related to time references within the news content (Shen et al., 2024; Mu et al., 2023). Modifications to dates, times, or the sequence of events can significantly impact the interpretation of the news. (6) Paraphrasing Degree (PD) evaluates the extent to which the content has been rephrased or paraphrased from the original text, which may obscure the original meaning or introduce ambiguity.
We employ LLMs to automate the evaluation process of the FUSE-EVAL metrics. Each of these dimensions is scored by GPT-4o-mini on a scale from 1 to 10, where a score of 1 indicates minimal or no deviation from the original news, and a score of 10 signifies a significant or complete deviation along that dimension. By assigning numerical scores, we can quantitatively compare the extent of changes across different aspects of the news content. As shown in the Figure 3, FUSE-EVAL demonstrates cumulative deviations (Pröllochs and Feuerriegel, 2023) during fake news evolution, confirming its effectiveness. To evaluate the overall deviation, we calculate the Total Deviation (TD) for each agent at each time step by averaging the scores:
(5) $\text{TD}_{i}^{t}=\frac{1}{6}\sum_{d=1}^{6}D_{i,d}^{t},$
where $D_{i,d}^{t}$ is the score of dimension $d$ for agent $i$ at time $t$. The detailed evaluation process are provided in Appendix B.
4.2. Statistical Deviation Metrics
The Statistical Deviation Metrics are derived from the Total Deviation scores and provide insights into the overall trends and patterns of news evolution within the network. We then analyze several statistical metrics based on the TD. The $\Delta$Deviation represents the difference in Average Deviation between the final and initial day of the simulation, highlighting the overall growth in deviation throughout the simulation period. The Average Deviation is the mean of TD across all agents at each time step, providing insight into the general trend of news evolution within the network. The Deviation Variance measures the statistical variance of TD among agents, indicating the degree of dispersion and how uniformly the news content deviates across the network. The Final Deviation is the average TD at the final simulation time step $t$, representing the cumulative effect of news evolution over the entire simulation duration. The Maximum Deviation and Minimum Deviation refer to the highest and lowest average TD observed during the simulation, respectively. These metrics indicate the points at which the news content has deviated most and least significantly from the original. The Peak Deviation Time indicates the percentage of the total simulation time taken to reach the Peak Deviation Rate. It provides insight into how quickly the maximum deviation occurs relative to the entire simulation duration. The Half Deviation Time is the time step $t_{0.5}$ at which the average TD reaches half of the Max Deviation, indicating the speed at which significant deviation occurs within the network.
4.3. Implementation Details
Our simulation framework was developed using Python scripts, leveraging various libraries to model the agents and their environment effectively. The LLM used is gpt-4o-mini, accessed via OpenAI API calls. When creating the network structure, we used the Python library networkx to construct different social network structures. The simulation includes 40 agents, whose traits were based on the Big Five personality dimensions commonly used in psychology (Barrick and Mount, 1991). Each agent was assigned scores on these traits to introduce variability in behaviors and interactions within the simulation. For further details, please refer to our code. We ensured that our framework is adaptable to different language models and have provided details on API costs and compatibility with other models in Appendix E and F.
Table 1. Comparative analysis of fake news evolution across different settings, including variations in topics, social networks, spread traits, and intervention strategies. Upward or downward arrows represent better control of fake news evolution.
Comparison Factors Setting $\Delta$ Deviation $\downarrow$
Average
Deviation $\downarrow$
Deviation
Variance $\downarrow$
Max
Deviation $\downarrow$
Min Deviation
Final
Deviation $\downarrow$
Peak Deviation
Time $\uparrow$
Half $\Delta$ Deviation
Time $\uparrow$
Topic Politics 3.148 6.594 0.511 7.440 3.442 6.590 0.133 0.033
Science 1.446 3.533 0.207 4.236 2.026 3.472 0.767 0.033
Scale-Free 2.631 4.287 0.725 5.652 1.492 4.955 0.767 0.167
High-Clustering 4.313 6.193 1.027 7.030 2.348 6.661 0.500 0.033
Emotional Spread 1.688 4.182 0.456 5.105 2.008 4.303 0.333 0.067
Super Spread 2.920 4.434 0.672 5.613 2.054 5.067 0.700 0.100
Vigilant 1.945 4.081 0.446 5.021 2.485 4.593 0.400 0.133
Intervention 1.384 4.207 0.476 5.302 1.841 4.559 0.200 0.067
Figure 3. (a) The FUSE-EVAL scores show cumulative deviations of the information over time. (b) Ablation study showing the effectiveness of hierarchical memory and propagation roles. (c) Impact of removing different agent types on fake news evolution. (d) Effectiveness of early-intervention, showing a clear reduction in deviation over time compared to the no-intervention condition.
5. Validation of the FUSE Framework
In this section, we demonstrate the effectiveness of our FUSE framework by showing that FUSE aligns with the influence of certain factors on fake news propagation, reproduces these patterns, and can also replicate real-world fake news.
5.1. Alignment with Real-World Patterns
5.1.1. Topic Comparison
We compared the evolution of fake news across five topics: politics, science, finance, terrorism, and urban legends. Table 1 shows that political fake news spreads the fastest, with average deviation peaking rapidly within four days and remaining high. Terrorism follows a similar pattern. In contrast, financial and science-related fake news spread more slowly, with science-related misinformation evolving the slowest and average deviation remaining consistently low. Table 1 shows the final deviation for political news is approximately 90% higher than that of science news. These results indicate that political fake news is more prone to rapid distortion and widespread belief, while science-related misinformation spreads more cautiously, consistent with prior research (Lazer et al., 2018b).
5.1.2. Social Network Comparison
Using a terrorism topic, we compared fake news evolution across different social network structures: random, scale-free, and high-clustering networks. Table 1 shows that high-clustering networks exhibit the fastest and most extensive spread of misinformation, with deviation peaking rapidly and remaining high. This indicates that tightly connected communities are particularly susceptible to rapid belief distortion, aligning with the ”echo chamber” effect (Cinelli et al., 2021). Random networks show the slowest evolution of fake news, with lower variance suggesting individuals are less likely to form polarized beliefs. Scale-free networks are intermediate. Peak deviation time is the longest in random networks and shortest in high-clustering networks, illustrating that clustering accelerates fake news evolution, consistent with prior research (Lind et al., 2007; Trpevski et al., 2010).
5.1.3. Spread Type Comparison
Using a terrorism topic, we compared fake news evolution across three spread types: normal spread, emotional spread, and super spread. Normal spread follows standard settings, super spread was assigned to high-degree nodes, and emotional spread involves participants using heightened emotional language. Table 1 shows that super spread leads to the highest level of misinformation, due to highly connected influencers amplifying dissemination. The emotional spread has a moderate effect, while normal spread shows the slowest evolution. Super spread reaches peak deviation fastest, followed by emotional spread; normal spread takes the longest. These results indicate that super spread and emotional spread drive faster and more significant misinformation evolution.
5.1.4. Personality Traits Comparison
Using a terrorism topic, we compared the impact of personality traits on fake news evolution. Based on the Big Five personality traits (Barrick and Mount, 1991), individuals with high agreeableness and neuroticism (Impressionable) are more likely to believe rumors, while those with low levels (Vigilant) are less susceptible. Table 1 show that impressionable agents are more prone to accepting and spreading misinformation. Vigilant agents exhibit stronger resistance, with beliefs remaining relatively stable. These results align with previous studies and demonstrate our framework’s effectiveness in simulating the influence of personality traits on fake news spread.
5.2. Alignment with Real-World Fake News
We conducted experiments across various topics and found that the fake news evolved by the FUSE framework closely corresponds to real-world fake news. As shown in Figure 4, the news about ”Trump being attacked” starts as true, evolves into partially false, and eventually becomes entirely fake. The agent, as a commentator, often adds its own views, while its neighboring verifiers and spreaders act according to their roles. Additionally, our framework generates fake news such as ”Trump was not attacked, it’s a dramatic effect,” which is also a widely circulated piece of fake news in the real world [^2] https://x.com/cwebbonline/status/1814708054916784594333https://x.com/EndWokeness/status/1813898763100176484. From a quantitative analysis perspective, for each topic, 73% of fake news is recovered by our framework. The detailed case study and analysis results are provided in the Appendix C.
6. ANALYSIS AND DISCUSSION
6.1. Ablation Study
6.1.1. The Impact of Hierarchical Memory and Propagation-Role.
As shown in Figure 3(b), the full FUSE framework demonstrates a clear accumulation effect in deviation, indicating its effectiveness in simulating fake news evolution. After removing hierarchical memory, the deviation significantly drops, with a 39.8% reduction throughout the simulation. The absence of the accumulation effect indicates the simulation fails (Pröllochs and Feuerriegel, 2023). This highlights the critical role hierarchical memory plays in capturing the depth and persistence of belief distortion, as it enables agents to store and reflect on both short-term and long-term information. Similarly, removing the propagation role further decreases the deviation. This underscores the importance of agent roles (such as spreader, commentator, verifier, and bystander) in shaping the propagation and evolution of information. Without these roles, the agents behave more uniformly, and the accumulation effect of deviation disappears, meaning that the news does not evolve.
6.1.2. The Impact of Propagation Role Types.
After the first ablation study showed that removing the propagation role leads to simulation failure, we conducted a second study to assess the impact of different role types (spreader, commentator, verifier, and bystander) on fake news evolution. As shown in Figure 3(c), removing the commentator agent had the most significant impact, with a sharp drop in average deviation, leading to simulation failure. This confirms the central role of the commentator in amplifying and disseminating misinformation. In contrast, removing the spreader slightly reduced the deviation, as spreaders do not provide explanations or opinions, which are key drivers of news distortion but may still contribute to its evolution. Interestingly, removing the verifier increased deviation, as verifiers play a crucial role in checking the accuracy of information. Without verifiers, the system becomes more susceptible to misinformation, leading to a greater spread of fake news. Finally, removing the bystander had the least impact, as bystanders passively observe without actively participating in the information spread.
Overall, these ablation studies demonstrate the effectiveness of FUSE. Hierarchical memory, propagation roles, and specific agent types play important roles in shaping the evolution of fake news within the FUSE framework.
6.2. Fake News Intervention
Figure 4. A Case of FUSE: A true news gradually evolves into partially false and eventually entirely fake news over time.
6.2.1. Intervention Strategy.
Based on the previous experimental results, we chose to introduce the official agent at high-degree nodes. Table 1 presents the comparative experimental results between the no-intervention and intervention conditions. Since fake news evolution reached a high point on the sixth day, we implemented the intervention strategy, where the official agent released a statement. The deviation under the intervention condition decreased by 37.8% compared to the no-intervention condition. This early intervention effectively blocked the initial spread of fake news, preventing a large number of agents from believing in the misinformation. However, as the news continued to propagate, the effectiveness of the intervention gradually weakened. By the 12th day, although the deviation under the intervention condition was still lower than that under the no-intervention condition, the gap narrowed to 22.3%, suggesting that the impact of a single intervention diminishes over time, as agents continue to interact and may revert to their previous beliefs or encounter conflicting information. Around the 16th day, we conducted a second intervention, which again reduced the deviation by 31.8% compared to the no-intervention condition. This further demonstrates the importance of continuous and timely fact-checks to suppress the spread of fake news. Compared to the no-intervention condition, the final deviation under the intervention strategy was reduced by approximately 28.6%  and the variance in deviation was also reduced by 61.8%. Moreover, the peak deviation time under the intervention strategy occurred approximately 0.56-time units earlier than in the no-intervention case.
Overall, the intervention strategy consistently maintained a lower average deviation than the no-intervention condition. The results highlight the importance of not only introducing early interventions but also maintaining regular interventions over a longer period to effectively combat the continuous spread and evolution of fake news.
6.3. Factors in Fake News Evolution
Figure 5. (a) The contribution percentages of factors in FUSE-EVAL to fake news evolution. (b) Comparison of the contributions of these factors across different topics, with Politics and Terrorism showing balanced contributions, while Science relies more on NII and less on TS and STS.
The analysis of experimental results and charts indicates varying contributions of different factors to fake news evolution. The Figure 5 (a) shows that PS contributes the most (22.3%), suggesting that altering reporting angles or distorting original information is the key driver of fake news evolution. NII follows with 18%, highlighting its significant role in this process. SS an STS contribute 17.2% and 17.5%, respectively, while TS has the smallest impact at 11%. The Figure 5 (b) reveals topic-specific patterns. Political and terrorism-related fake news evolves across multiple dimensions, especially new information, perspective, and sentiment shifts. In contrast, science-related fake news is driven mainly by new information, with less influence from temporal or style shifts. Urban legends and finance topics also heavily rely on perspective shifts and new information. In summary, PS and NII are the main drivers of fake news evolution, with time-related changes having the least impact. Understanding these patterns can help in developing targeted strategies to detect and mitigate misinformation.
7. CONCLUSION
We introduced the FUSE framework, a simulation environment that models the evolution of true news into fake news using agent-based modeling and LLMs. By implementing the FUSE-EVAL evaluation framework, which quantifies content deviation across six dimensions, we systematically analyzed how misinformation propagates and transforms within social networks. Our experiments demonstrated alignment with real-world observations across various settings, reinforcing existing theories on misinformation spread, such as the rapid distortion of political news, the influence of network clustering, the impact of super spreaders and emotional content, and the role of individual personality traits in susceptibility to fake news. By automating the evaluation process with LLMs like GPT-4o-mini, our framework achieves scalable and consistent analysis. By enhancing our understanding of misinformation spread, our framework supports the development of a more responsible and trustworthy web environment.
Appendix A Prompt Set

### v2

@@ 3.3.2 Role-Specific Behaviors
$f_{role}=f_{r_{i}}(S_{i}^{t-1},\{S_{j}^{t-1}|a_{j}\in\mathcal{N}_{i}\},\mathcal{P}_{i}).$ (2)
@@
For different roles in our model, spreaders may combine and amplify sensational aspects of the news, commentators may add personal opinions, verifiers may check news before sharing, and bystanders may retain their previous news content unless significantly influenced Sun et al. (2023).
3.3.3 Memory and Reflection
@@
In our simulation, agents engage with their neighbors daily, leading to updated versions of the news. Given the volume of interactions, we implement a hierarchical memory system comprising short-term memory (STM) $M_{i}^{S}$ for recent interactions and long-term memory (LTM) $M_{i}^{L}$ for accumulated knowledge. After interactions, agents reflect and update the news through a memory function:
$M_{i}^{L,t}=g(f_{L}(M_{i}^{L,t-1}),f_{S}(M_{i}^{S,t})),$ (3)
@@
where $g(\cdot)$ integrates new information into LTM, enabling agents to exhibit dynamic behaviors such as gradually changing their opinion on a topic or reinforcing existing opinions.
3.3.4 Decision-Making Process
@@
In our FUSE framework, each agent’s opinion evolves through a reasoning process influenced by their role, persona, and interactions. Agents reflect on their news content after daily interactions and memory updates, leading to gradual opinion changes. The decision-making process for agent $a_{i}$ at time $t$ is modeled as:
$S_{i}^{t}=f_{dm}(S_{i}^{t-1},m_{i}^{L,t-1},r_{i},\mathcal{P}_{i}).$ (4)
@@
This function captures how agents integrate new information with their existing opinions, considering their role in the decision-making process. For example, the reasoning of spreaders may lead to greater changes in $S_{i}^{t}$, commentators add subjective nuances, verifiers aim to correct inaccuracies, and bystanders typically make minimal changes.
3.4 News Evolution Simulator
The News Evolution Simulator (NES) provides the environment where news content evolves over time through agent interactions within a social network structure $\mathcal{G}=(\mathcal{A},\mathcal{E})$. This module enables studying how true news transforms into fake news through agent behaviors and social interactions.
NES models various network topologies to reflect different social dynamics: random networks with randomly formed edges between agents $a_{i}\in\mathcal{A}$, simulating loosely connected environments; scale-free networks with hub agents acting as “super-spreaders”; and high-clustering networks forming tightly-knit communities that mirror real-world social circles Nekovee et al. (2007); Moreno et al. (2004). As outlined in Appendix I, the network structure $\mathcal{G}$ determines daily agent interactions, influencing news content’s evolution patterns. The overall algorithm is presented in Appendix A.
3.4.1 Intervention Mechanisms
A key feature of NES is its ability to simulate interventions to counter fake news evolution. When the deviation between current news content $S_{i}^{t}$ and original news $S_{0}$ exceeds a predefined threshold, an official agent is introduced to provide verified information and correct misinformation.
The intervention process starts with continuously monitoring the deviation between each agent’s news content and the original news. Once the deviation exceeds a critical threshold, the official agent is triggered to take action. This agent issues official announcements based on reliable sources, targeting agents most likely to propagate or exacerbate misinformation.
The prompts for all functions mentioned in § 3 can be found in Appendix B.
4 FUSE-EVAL: News Evolution Analysis
To systematically measure how true news evolves into fake news within our simulation, we propose a comprehensive evaluation framework named FUSE-EVAL. This framework consists of two sets of metrics: Content Deviation Metrics and Statistical Deviation Metrics, which together provide a detailed understanding of how fake news evolves within the simulated environment.
4.1 Content Deviation Metrics
The Content Deviation Metrics assess the deviation of the news content across multiple dimensions by quantifying changes in specific aspects of the news. FUSE-EVAL evaluates the news content based on six core dimensions:
(1) Sentiment Shift (SS) measures the change in emotional tone between the original news content and its evolved version Lu et al. (2022); Ma et al. (2021). Sentiment plays a crucial role in how information is perceived and shared, with shifts indicating potential bias or emotional manipulation.
(2) New Information Introduced (NII) assesses the extent to which additional information, not present in the original news, has been incorporated Wang et al. (2017). Introducing new facts or claims can significantly alter the original message, potentially leading to misinformation.
(3) Certainty Shift (CS) evaluates changes in the level of confidence or assertiveness expressed in the news content Krafft et al. (2019); Kim and Yoon (2022). Shifts from definitive to speculative language can influence the perceived credibility of information.
(4) Stylistic Shift (STS) examines changes in writing style, tone, and linguistic features Wu et al. (2024). Alterations in style can affect readability and audience engagement through formality and sentence complexity changes.
(5) Temporal Shift (TS) measures changes related to time references within the news content Shen et al. (2024); Mu et al. (2023). Modifying dates, times, or event sequences can significantly impact news interpretation.
(6) Paraphrasing Degree (PD) evaluates the extent to which the content has been rephrased from the original text, which may obscure meaning or introduce ambiguity.
We employ GPT-4o-mini to automate FUSE-EVAL evaluation, scoring each dimension from 1 (minimal deviation) to 10 (significant deviation).
As shown in Figure 3 (a), FUSE-EVAL demonstrates cumulative deviations Pröllochs and Feuerriegel (2023) during fake news evolution, confirming its effectiveness. To evaluate the overall deviation, the Total Deviation (TD) for each agent at each time step $t$ is calculated as:
$\text{TD}_{i}^{t}=\frac{1}{6}\sum_{d=1}^{6}D_{i,d}^{t},$ (5)
where $D_{i,d}^{t}$ is the score of dimension $d$ for agent $i$ at time $t$. The detailed evaluation process is provided in Appendix D.
Figure 3: (a) The FUSE-EVAL scores show cumulative information deviations over time. (b) A Case of FUSE: True news gradually evolves into partially false and eventually entirely fake news over time.
4.2 Statistical Deviation Metrics
The Statistical Deviation Metrics, derived from Total Deviation (TD) scores, provide insights into the overall patterns of news evolution within the network. We analyze several key metrics:
$\bullet$ The $\Delta$Deviation represents the difference in Average Deviation between the final and initial simulation day, indicating overall deviation growth.
$\bullet$ The Average Deviation is the mean of TD across all agents at each time step, showing the general trend of news evolution within the network.
$\bullet$ The Deviation Variance measures the statistical variance of TD among agents, measuring how uniformly content deviates across the network.
$\bullet$ The Final Deviation is the average TD at the finaltime step $t$, representing the cumulative effect.
$\bullet$ The Maximum Deviation and Minimum Deviation refer to the highest and lowest average TD observed, showing the extremes of news deviation.
$\bullet$ The Peak Deviation Time indicates the percentage of simulation time taken to reach Peak Deviation Rate, showing the speed of maximum deviation occurrence.
$\bullet$ The Half Deviation Time is the time step $t_{0.5}$ when average TD reaches half of Max Deviation, indicating the rate of significant deviation.
4.3 Implementation Details
Our framework uses GPT-4o-mini as the primary LLM and the simulation comprises 40 agents. Additional implementation details, including agent personality traits and programming environment, are provided in Appendix C. At the same time, API costs and compatibility with other models can be found in Appendix G and Appendix H.
Comparison Factor Setting
[…]
Average
[…]
Deviation
[…]
Max
[…]
Min
Deviation
Final
[…]
Peak Deviation
Time $\uparrow$
Half $\Delta$ Deviation
[…]
Topic Politics 3.148 6.594 0.511 7.440 3.442 6.590 0.133 0.033
Science 1.446 3.533 0.207 4.236 2.026 3.472 0.767 0.033
Scale-Free 2.631 4.287 0.725 5.652 1.492 4.955 0.767 0.167
High-Clustering 4.313 6.193 1.027 7.030 2.348 6.661 0.500 0.033
Emotional Spread 1.688 4.182 0.456 5.105 2.008 4.303 0.333 0.067
Super Spread 2.920 4.434 0.672 5.613 2.054 5.067 0.700 0.100
Vigilant 1.945 4.081 0.446 5.021 2.485 4.593 0.400 0.133
Intervention 1.384 4.207 0.476 5.302 1.841 4.559 0.200 0.067
Table 1: Comparative analysis of fake news evolution across different settings, including variations in topics, social networks, spread traits, and intervention strategies. $\uparrow$ or $\downarrow$ arrows represent better control of fake news evolution. Bold numbers indicate statistically significant improvements over baseline models (t-test with p-value$<$0.01).
5 Validation of the FUSE Framework
In this section, we demonstrate FUSE’s effectiveness by validating its alignment with known fake news propagation patterns and its ability to reproduce real-world fake news.
5.1 Alignment with Real-World Patterns
Topic Comparison
We analyzed fake news evolution across five topics: politics, science, finance, terrorism, and urban legends. As shown in Table 1 and Appendix F, political fake news exhibits the fastest spread, with average deviation peaking within four days, followed by terrorism-related content. Science and financial news evolve more slowly, showing the lowest average deviation. Table 1 shows the final deviation for political news is approximately 90% higher than that of science news. These results indicate that political fake news is more prone to rapid distortion and widespread belief, while science-related misinformation spreads more cautiously, aligned with prior research Lazer et al. (2018a). We collected 120 pieces of true news across five topics. All news is published after the training cutoff date of GPT-4o-mini. The results were consistent, and the dataset will be publicly available.
Social Network Comparison
We analyzed fake news evolution across three network structures (random, scale-free, and high-clustering) using a terrorism topic. Table 1 shows that high-clustering networks lead to the fastest and most extensive fake news spread, with deviation peaking rapidly and remaining high. This indicates that tightly connected communities are particularly susceptible to rapid belief distortion, aligning with the “echo chamber” effect Cinelli et al. (2021). Random networks show the slowest evolution of fake news with lower variance, while scale-free networks exhibit intermediate behavior. Peak deviation time is the longest in random networks and shortest in high-clustering networks, illustrating that clustering accelerates fake news evolution, consistent with prior research Lind et al. (2007); Trpevski et al. (2010).
Spread Type Comparison
We analyzed three spread types (normal, emotional, and super spread) using a terrorism topic. Super spread, assigned to high-degree nodes, leads to the highest misinformation level due to influencer amplification. Emotional spread, characterized by heightened emotional language, shows moderate effects, while normal spread exhibits the slowest evolution. As shown in Table 1, peak deviation time is shortest in super spread, followed by emotional spread, demonstrating their accelerating effect on misinformation evolution, aligned with prior research Sun et al. (2023).
Personality Traits Comparison
Using a terrorism topic, we compared the impact of personality traits on fake news evolution. Based on the Big Five personality traits Barrick and Mount (1991), we compared agents with high agreeableness and neuroticism (Impressionable) versus low levels (Vigilant). Table 1 shows that Impressionable agents are more prone to accepting and spreading misinformation. In contrast, Vigilant agents maintain more stable beliefs, aligning with previous studies on personality influence in fake news spread Mirzabeigi et al. (2023).
5.2 Alignment with Real-World Fake News
[…]
6 Analysis and Discussion
Figure 4: (a) Ablation study showing the effectiveness of hierarchical memory and propagation roles. (b) Impact of removing different agent types on fake news evolution. (c) Effectiveness of early intervention, showing an apparent reduction in deviation over time compared to the no-intervention condition.
6.1 Ablation Study
The Impact of Hierarchical Memory and Propagation-Role.
As shown in Figure 4 (a), the complete FUSE framework demonstrates apparent deviation accumulation, indicating its effectiveness in simulating fake news evolution. After removing hierarchical memory, the deviation significantly drops, with a 39.8% reduction throughout the simulation, indicating the simulation fails Pröllochs and Feuerriegel (2023). This highlights memory’s crucial role in capturing persistent belief distortion through short-term and long-term information processing. Similarly, removing propagation roles leads to further deviation decrease, emphasizing how distinct agent roles (spreader, commentator, verifier, and bystander) shape information evolution. Without these roles, the agents behave more uniformly, and the accumulation effect of deviation disappears, meaning that the news does not evolve.
The Impact of Propagation Role Types.
Following our first ablation study showing that removing propagation roles leads to simulation failure, we conducted a detailed analysis of different agent roles’ impact on fake news evolution. As shown in Figure 4 (b), removing commentators caused the most significant drop in average deviation, confirming their crucial role in false news spread through opinion addition and interpretation. Removing spreaders had a relatively minimal impact as they lack opinion-adding capabilities, though they still contribute to information dissemination.
Removing verifiers increased overall deviation, demonstrating their important role in maintaining information accuracy through fact-checking. Without verifiers, the system became more susceptible to misinformation spread. Bystander removal showed the least effect, consistent with their passive observational role in the network.
These findings, combined with our previous ablation results on hierarchical memory and propagation roles, validate FUSE’s effectiveness and demonstrate how different components contribute to simulating fake news evolution.
6.2 Fake News Intervention Strategy
Based on previous results, we implemented interventions through an official agent at high-degree nodes. As shown in Table 1 and Figure 4 (c), when fake news evolution peaked on the sixth day, our first intervention reduced deviation by 37.8% compared to no-intervention. Although this effect gradually weakened, with the gap narrowing to 22.3% by day 12 as agents continued to interact and potentially revert to previous beliefs, a second intervention on day 16 achieved a 31.8% reduction in deviation. The intervention strategy demonstrated several significant improvements over the no-intervention condition: the final deviation decreased by 28.6%, the deviation variance reduced by 61.8%, and the peak deviation occurred 0.56 time units earlier. Throughout the simulation, the intervention strategy consistently maintained lower average deviation levels. These results emphasize that effective fake news mitigation requires both early and regular interventions to combat the continuous evolution of fake news.
6.3 Factors in Fake News Evolution
Figure 5: (a) The contribution percentages of factors in FUSE-EVAL to fake news evolution. (b) Comparison of the contributions of these factors across different topics, with Politics and Terrorism showing balanced contributions, while Science relies more on NII and less on TS and STS.
The analysis of experimental results and charts indicates varying contributions of different factors to fake news evolution. Figure 5 (a) shows that PD contributes the most (22.3%), suggesting that altering reporting angles or distorting original information is the key driver of fake news evolution. NII follows with 18%, highlighting its significant role in this process. SS an STS contribute 17.2% and 17.5%, respectively, while TS has the most negligible impact at 11%. Figure 5 (b) reveals topic-specific patterns. Political and terrorism-related fake news evolves across multiple dimensions, especially new information, perspective, and sentiment shifts. In contrast, science-related fake news is driven mainly by new information, with less influence from temporal or style shifts. Urban legends and finance topics rely heavily on perspective shifts and new information. In summary, PD and NII are the main drivers of fake news evolution, with time-related changes having the least impact. Understanding these patterns can help develope targeted strategies to detect and mitigate fake news.
7 Conclusion
We presented FUSE, a framework that simulates the evolution of true news into fake news using LLM-based agents. Through our FUSE-EVAL framework, which measures content deviation across six dimensions, we analyzed fake news evolution patterns in social networks. Our experiments validated several established theories, including the accelerated spread of political fake news, effects of network clustering, impacts of super spreaders and emotional content, and the role of personality traits in fake news susceptibility. Using LLMs for automated evaluation enables scalable analysis, contributing to understand the fake news dynamics.
Limitations
Despite the advancements presented by FUSE, our study faces three primary limitations:
Data Availability: Currently, there is a lack of comprehensive datasets that capture the dynamic process of fake news evolving from true information. Most existing datasets focus on static instances of misinformation or their immediate spread, which restricts our ability to fully validate FUSE across diverse real-world scenarios.
Complex Social Factors: Our current framework focuses on key social dynamics and individual personality traits in fake news evolution, without explicitly modeling broader factors such as political agendas, ideological bias, or crisis-driven contexts. These complex social factors can influence how true news is distorted in real-world settings. Nevertheless, the modular design of FUSE allows future extensions to incorporate such context for more comprehensive simulations.
Evaluation Methodology: Our evaluation framework, FUSE-EVAL, relies on specific dimensions such as Sentiment Shift and New Information Introduced to measure deviations in news content. However, these metrics may not cover all aspects of fake news evolution, potentially missing subtle nuances in misinformation dynamics. Additionally, the dependence on LLMs for simulation and evaluation may introduce inherent biases, affecting the accuracy of our assessments.
Appendix A The Overall Algorithm
@@
Appendix B Prompt Set